OpenStudy (anonymous):

find the volume of the solid generated by revolving the region bounded by the graphs of the equations about the -axis. y= 3/(x+1) y=0 x=0 x=8

OpenStudy (amistre64):

your missing part of the question

OpenStudy (anonymous):

Am I? let me check.

OpenStudy (jhannybean):

Does your question tell you which axis it wants the function revolved around? Or can it be either one?

OpenStudy (anonymous):

Thats all they gave me. I think i know how to set it up though. I believe its the x axis

OpenStudy (amistre64):

and what are your thoughts on how to approach it?

OpenStudy (anonymous):

I believe I know how to set this problem up, you use the disk method and the boundaries are [0,8]

OpenStudy (amistre64):

good, sound ideas

OpenStudy (anonymous):

I just forgot some basic steps

OpenStudy (amistre64):

such as?

OpenStudy (amistre64):



OpenStudy (anonymous):

\[pie \int\limits_{0}^{8} [3/(x+1]^2\]

OpenStudy (amistre64):

teh area of a circle is pi r^2, and r is defined as y for a given value of x yep

OpenStudy (anonymous):

Why question is do i square the denominator or just the 3

OpenStudy (anonymous):

*one

OpenStudy (amistre64):

pi is a greek letter pie is an edible disc :)

OpenStudy (amistre64):

y^2, what does y=?

OpenStudy (anonymous):

I know, sorry my bad

OpenStudy (anonymous):

I think i am finding the area under y=3/x+1 bounded by y=0 x=0 x=8

OpenStudy (anonymous):

So i use the disk method since there inst a . Right?

OpenStudy (amistre64):

\[\int \pi r^2\] \[\int \pi (y)^2~dx\] \[\pi\int (\frac{3}{x+1})^2~dx\] \[\pi\int \frac{3}{x+1}*\frac{3}{x+1}~dx\] \[\pi\int \frac{9}{(x+1)^2}~dx\]

OpenStudy (jhannybean):

Not quite. You're finding the area of the revolved function... which is your y\(^2\)

OpenStudy (jhannybean):

Basically your volume.

OpenStudy (anonymous):

Thats what i first had, but i forgot how to integrate that. Do i have to use ln?

OpenStudy (jhannybean):

\[V=\int A(x)dx\] Remember? :)

OpenStudy (amistre64):

try it, whats the derivative of ln((x+1)^2) ??

OpenStudy (anonymous):

2(x+1)/(x+10^2?

OpenStudy (amistre64):

2x/(x+1)^2 so ln aint gonna cut it, but its simler than that

OpenStudy (amistre64):

\[\pi\int \frac{9}{(x+1)^2}~dx\] \[\pi\int 9(x+1)^{-2}~dx\] its just a power rule

OpenStudy (anonymous):

okay , sorry i had a brain fart. Let me try to work it out by myself first

OpenStudy (amistre64):

whats the derivative of (x+1)^(-1) ??

OpenStudy (anonymous):

Derivative or anti derivative

OpenStudy (amistre64):

well im jsut trying to get you to recall the power rule, and since we drop down to -2, it must come from -1

OpenStudy (anonymous):

I remember how to do it. I will try to solve it from here, but can u just check my answer?

OpenStudy (amistre64):

im feeling pretty lucid, so sure :)

OpenStudy (anonymous):

8 pi?

OpenStudy (amistre64):

good job

OpenStudy (anonymous):

Thanks for your help, I over complicate things some times

OpenStudy (amistre64):

lol, i used to make up my own problems during a test, and solve them ... got the right answer, but not for the actual problem on the test :)

OpenStudy (amistre64):

good luck :)